Using the given equation find the missing coordinates of the points and then find the slope of the line for each equation:
y=3.5x+4; A(_;5) B(10;_)
7x-5y=12; A(_;9) B(4;_)
y=−3/7x+3; A(_;6) B(9;_)
y=0.3x+1; A(_;9), B(5;_)

Answer by macston(5194)

.
y=3.5x+4; A(_,5) B(10,_)
.
Find point A: Substitute y value given, solve for x:
.
5=3.5x+4
1=3.5x
1/3.5=x
0.2857=x
.
Point A=(0.2857,5)
.
Find point B: Substitute x value given, solve for y:
.
y=3.5(10)+4
y=35+4
y=39
.
Point B=(10,39)
.
Slope:
In slope intercept form y=mx+b, m=slope
.
Slope=m=3.5
.
The others are done the same way.